[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 06, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 6 08:36:59 UTC 2016


HSU Shift Log for Wed Jul 06 2016 04:36:53 GMT-0400 (EDT)

Shift: A

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: The TWO text products are not updating regularly in  
NWX. For example, the latest TWOAT is from 18Z Tuesday, but the TWOEP is up  
to date. Someone should check the LDM feed into $TEXT_DATA and $TEXT_DATA2.

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160706/a197c149/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list